OpenStudy (anonymous):

-1/2 + 3/4 In simplest form Medals will be awarded

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

that is the correct answer

OpenStudy (anonymous):

\[\large \frac{3}{4}-\frac{1}{2}=\frac{3}{4}-\frac{2}{4}\] is as start

OpenStudy (anonymous):

i have a link to help you every time for questions here http://www.helpwithfractions.com/fraction-calculator/

OpenStudy (isaiah.feynman):